OpenStudy (anonymous):

I need help with Rational Root Theorem and Descartes Rule of Signs. What exactly is the Rational Root Theorem and Descartes Rule of Signs? and how do I answer these three questions?? 1. What are the possible rational zeros of f(x) = x^4 + 6x^3 - 3x^2 + 17x - 15? 2. What are the possible rational zeros of f(x) = 3x^4 + x^3 - 13x^2 - 2x + 9? 4. What are the possible number of positive real, negative real, and complex zeros of f(x) = 4x3 + x2 + 10x - 14?

OpenStudy (anonymous):

@phi Can you walk me through how to do these problems please?

OpenStudy (phi):

Here is Descartes http://www.purplemath.com/modules/drofsign.htm

OpenStudy (anonymous):

okay I have a better understanding of Descartes

OpenStudy (phi):

For questions 1 and 2, you use the rational root theorem This video goes into it. http://www.youtube.com/watch?v=YMyv9-9VXw4

OpenStudy (anonymous):

Thanks be right back, going to watch it.

OpenStudy (anonymous):

@phi After watching I solved #1 and came of with the numbers 1,6,-3,17,-15 as possible rational zeros.

OpenStudy (anonymous):

I think I did it correctly... 1. What are the possible rational zeros of f(x) = x^4 + 6x^3 - 3x^2 + 17x - 15? ± 1, ± 3, ± 5, ± 15 ± 1, ± 1 over 3, ± 1 over 5, ± 1 over 15 ± 1, ± 3, ± 6, ± 15, ± 17 <-- is the answer I would choose. ± 1, ± 1 over 3, ± 1 over 6, ± 1 over 15, ± 1 over 17

OpenStudy (phi):

The Rule is "factors of the constant term" (the number with no x next to it) divided by the "factors of the highest-order term" (factor of the x to the biggest exponent) in this case you look at 1x^4 and -15, and see you want 15 and 1 it will be the factors of 15 divided by the factors of 1. factors of 15 are 1,3,5,15 they can be ± , so we could list them as -1,-3,-5,-15,1,3,5,15

OpenStudy (phi):

Notice you do not use *all* the coefficients (which is what you did). You pick that very last number, and the very first coefficient (if the polynomial is written in standard form)

OpenStudy (phi):

For Question 2) f(x) = 3x^4 + x^3 - 13x^2 - 2x + 9 what two numbers do you use?

OpenStudy (anonymous):

3 and 9?

OpenStudy (phi):

yes. which goes "up top" ?

OpenStudy (anonymous):

1 and 3

OpenStudy (phi):

the factors of 9 (the last number in 3x^4 + x^3 - 13x^2 - 2x + 9) goes up top the factors of 3 (the first coefficient, from 3x^4 ) goes in the bottom \[ \frac{\text{ factors of 9}}{\text{factors of 3} }\]

OpenStudy (phi):

with ±

OpenStudy (anonymous):

yes \[\pm 1 \pm 3\] is that correct or would \[\pm 9\] be included

OpenStudy (phi):

the factors of 9 are 1,3,9 (the 3*3 =9 repeats the 3. we need to list it only once) the factors of 3 are 1,3 now find all combinations of 1,3, 9 divided by 1 and then by 3

OpenStudy (phi):

in other words, you want 1,3,9 divided by 1, which gives 1,3,9 then you want 1,3,9 divided by 3, which gives 1/3, 1,3 merge the list, and ignore duplicates what do you get ?

OpenStudy (anonymous):

± 1, ± 1/3, ± 3, ± 9

OpenStudy (phi):

that is the answer. Can you go over the process, and see how you get it ?

OpenStudy (anonymous):

Yes, Sir I have to keep really good notes cause my work is online and its harder to learn math. So to learn I've got to keep well written steps on how to find things.

OpenStudy (phi):

I don't see Question 3. Question 4 uses Descartes rule of signs.

OpenStudy (anonymous):

Question 3 was graphing a function. and I already knew how to do that... 4. What are the possible number of positive real, negative real, and complex zeros of f(x) = 4x3 + x2 + 10x - 14? A. Positive Real: 1 Negative Real: 0 Complex: 2 B. Positive Real: 1 Negative Real: 2 or 0 Complex: 2 or 0 C. Positive Real: 2 or 0 Negative Real: 1 Complex: 2 or 0 D. Positive Real: 2 or 0 Negative Real: 2 or 0 Complex: 1

OpenStudy (anonymous):

@phi

OpenStudy (phi):

See if this helps http://www.youtube.com/watch?v=MVX2vYlc24k

OpenStudy (anonymous):

ok